# Steiner Traveling Salesman Problem with Quantum Annealing

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; July 14 - 18, 2025; Málaga, ESisbn: 978-1-4503-XXXX-X/2018/06ccs: Applied computing Transportationccs: Theory of computation Quantum computation theory
@@ 1. Introduction
In a different vein, quantum computing has emerged as an innovative and promising approach to solving combinatorial optimization problems, particularly those classified as NP-hard. Thanks to quantum mechanical principles such as superposition and entanglement, quantum algorithms can explore many possible solutions simultaneously, increasing the potential to find optimal configurations in less time than traditional methods. For a comprehensive survey on quantum approaches used to solve routing problems with quantum algorithms, the reader is referred to (Ciacco et al., 2025b) and (Osaba et al., 2022).
Among the various paradigms of quantum algorithms, Quantum Annealing (QA) has attracted special interest as one of the most promising methods for efficiently solving complex combinatorial optimization problems. QA takes its name from an analogous physical process, annealing, which involves a material being heated and then cooling down slowly. This process allows the material to settle back into a state of low energy, ultimately reaching a low-energy location in a solution space. Additionally, quantum mechanical properties are used to effectively explore the possible states of the material.
@@
This study is the first to apply a quantum computing approach to solving the STSP. Specifically, we introduce the use of QA for this task, representing a pioneering effort in applying quantum methods to such a complex optimization problem. In particular, we analyze QA by using two different approaches: the D-Wave Quantum Processing Unit (QPU) and the D-Wave’s LeapBQM hybrid solver (Osaba et al., 2024). We provide a comprehensive evaluation of these two distinct quantum approaches to tackle the STSP. In addition to this, we introduce a network reduction method to enhance the QUBO representation of the problem, reducing the number of variables that would otherwise grow significantly when transforming the ILP formulation into QUBO. This formulation constitutes another significant contribution of this research.
The remaining of this paper is structured as follows. Section 2 presents the problem description, the mathematical formulation and the proposed preprocessing method for reducing arcs. Section 3 describes the proposed solution approach for the problem. Section 4 presents the computational tests and analysis of the results. Section 5 summarizes the concluding remarks and prospects for future studies.
@@ 2.1. Characteristics of the problem
Figure 1. Representation of the STSP network: the yellow circles represent the $V_{R}$ nodes, the green circle indicates the depot and the blue circles represent the steiner nodes. A possible route is highlighted in red, showing one of the potential solutions to the optimization problem.
@@
Graph representation of the STSP network. The graph consists of yellow, blue, and green circles connected by black and red directed edges. The yellow circles represent the $V_{R}$ nodes, the green circle is the depot, and the blue circles are the Steiner nodes. A possible route is highlighted in red, illustrating one potential solution to the optimization problem.
@@ 2.2. Mathematical Formulation
(7) $y_{k}^{t}\in\{0,1\},\quad\forall k\in A,1\leq t\leq|A|-1$
The objective function (1) minimizes the total travel cost across all time periods. It is expressed as the sum of the costs associated with each arc traversed, weighted by the respective cost. Constraint (2) ensures that the route starts at the depot. Constraints (3) ensure that in the first period, there are no other arcs traversed except for the one originating from the depot. Constraint (4) ensures that the sum of incoming arcs at each node must equal the sum of outgoing arcs. Constraints (5) enforce the requirement that all required vertices must be visited at least once. Constraints (6) ensure flow conservation between consecutive time periods for each node (except the depot). They state that the number of incoming arcs to a vertice in any given period must equal the number of outgoing arcs from that node in the next period, thereby maintaining a continuous flow through the network. Constraints (7) define the domain of the decision variable.
Table 1. Summary of the notations used in the model formulation.
Notation Description
$V$ Set of all vertices in the graph
$A$ Set of directed arcs in the graph
$V_{R}$ Set of the required nodes to visit
$\delta^{+}(i)$ Set of arcs outgoing from node $i$
$\delta^{-}(i)$ Set of arcs incoming to node $i$
$c_{k}$ Cost associated with traversing arc $k$
@@ 3. Solution approach
    QUBO formulation: we convert both the preprocessed and unpreprocessed versions of the STSP into QUBO formulations. The preprocessed version will be referred to as RQUBO, while the unprocessed version will be referred to as SQUBO. The QUBO formulation is solved using two different approaches:
@@ 4. Computational study
For the instances solved using Gurobi, we have use the Gurobi ver. 11 (Gurobi Optimization, LLC, 2024) and the implementation is done entirely in Python. Also, we impose a time_limit of 10 seconds. For SILP and RILP solved with Gurobi, we consider 16 instances with increasing sizes, ranging from 4 to 19 nodes. The analysis stops beyond this threshold, as no solution is obtained within the time limit.
@@
Experimental tests are conducted within a 100x100 unit square region. The set of nodes $V$ is randomly generated by assigning each node random Cartesian coordinates within this area. An additional node, denoted as the depot node 00, is generated with random coordinates within the same region. The subset of required nodes $V_{R}\subseteq V$ is selected comprising the first $70\%$ of the total vertices in $V$, rounded down to the nearest integer. Furthermore, each arc cost $c_{k}$ is assigned a random value between 20 and 50. To facilitate the replication of this study, all instances used are publicly accessible in (Ciacco et al., 2025a).
@@
  • Finally, the third section explores the performance of QA by employing the D-Wave solver.
@@ 4.1. Scalability of PMRA using Gurobi solver
We then transform both our SILP and RILP formulations into QUBO and evaluate the impact of the PMLA in reducing the problem size. Table 3 presents a comparison of the number of variables in the SQUBO and RQUBO formulations, highlighting the effectiveness of the PMLA in reducing the number of variables across different problem sizes. It consists of the column $|V|$, representing the cardinality of the set $|V|$, followed by n.var SQUBO and n.var RQUBO, which indicate the number of variables in the SQUBO and RQUBO formulations, respectively. Finally, the GAP column shows the percentage reduction achieved. Analyzing the results, we observe that the number of variables in the SQUBO formulation grows rapidly as $|V|$ increases. When the PMRA is used, the increase in the number of variables as the size of the set V grows is more contained. In fact, a maximum reduction of 57% and an average reduction of 48% are achieved.
@@ 5. Conclusions
We utilized both the D-Wave QPU and the D-Wave LeapBQM solver in our experiments. The LeapBQM consistently outperformed the QPU-based approach, both in terms of execution time and solution quality. While the QPU demonstrated effectiveness for very small problem instances, it was generally less efficient for larger problems. The LeapBQM provided high-quality solutions and showed better scalability, even for larger problem sizes. This makes the LeapBQM a more promising and practical tool for addressing real-world STSP challenges, offering superior performance across a range of problem sizes. Furthermore, in all the tests conducted across all methods, ILP solved with Gurobi, SA and both quantum approaches (QPU and LeapBQM), we observed that the PMRA consistently enhanced the solving performance. This method proved to be effective in improving the computational efficiency and the quality of the solutions, underlining its importance as a significant contribution to solving the STSP. In the future, we plan to extend our approach to larger logistics networks and more complex scenarios.
@@ Acknowledgments
Acknowledgments removed for anonymization purposes. During the preparation of this work, the authors used Microsoft Copilot to improve the language and readability of the manuscript. After using this tool/service, the authors have reviewed and edited the content as needed and take full responsibility for the content of the publication.

### v3

@@ Abstract.
journalyear: 2025copyright: ccconference: Genetic and Evolutionary Computation Conference; July 14–18, 2025; Malaga, Spainbooktitle: Genetic and Evolutionary Computation Conference (GECCO ’25 Companion), July 14–18, 2025, Malaga, Spaindoi: 10.1145/3712255.3734313isbn: 979-8-4007-1464-1/2025/07ccs: Applied computing Transportationccs: Theory of computation Quantum computation theory
@@ 1. Introduction
In a different vein, quantum computing has emerged as an innovative and promising approach to solving combinatorial optimization problems, particularly those classified as NP-hard. Thanks to quantum mechanical principles such as superposition and entanglement, quantum algorithms can explore many possible solutions simultaneously, increasing the potential to find optimal configurations in less time than traditional methods. For a comprehensive survey on quantum approaches used to solve routing problems with quantum algorithms, the reader is referred to (Ciacco et al., 2025a) and (Osaba et al., 2022).
Among the various paradigms of quantum algorithms, Quantum Annealing (QA) has attracted special interest as one of the most promising methods for efficiently solving complex combinatorial optimization problems. QA takes its name from an analogous physical process, annealing, which involves a material being heated and then cooling down slowly. This process allows the material to settle back into a state of low energy, ultimately reaching a low-energy location in a solution space. Additionally, quantum mechanical properties are used to explore the possible states of the material.
@@
This study is the first to apply a quantum computing approach to solving the STSP. Specifically, we introduce the use of QA for this task, representing a pioneering effort in applying quantum methods to such a complex optimization problem, such as STSP. In particular, we analyze QA by using two different approaches: the D-Wave Quantum Processing Unit (QPU) and the D-Wave’s LeapBQM hybrid solver (Osaba et al., 2024). We provide a comprehensive evaluation of these two distinct quantum approaches to tackle the STSP. In addition to this, we introduce a network reduction method to enhance the QUBO representation of the problem, reducing the number of variables that would otherwise grow significantly when transforming the ILP formulation into QUBO. This formulation constitutes another significant contribution of this research.
The remaining of this paper is structured as follows. Section 2 presents the problem description, the mathematical formulation and the proposed preprocessing method for reducing arcs. Section 3 describes the proposed solution approach for the problem. Section 4 presents the computational tests and analysis of the results. Section 5 summarizes the concluding remarks.
@@ 2.1. Characteristics of the problem
Figure 1. Representation of the STSP network: the yellow circles represent the $V_{R}$ nodes, the green circle indicates the depot and the blue circles represent the steiner nodes. The black edges represent the arcs of the graph, while a possible route is highlighted in red, showing one of the potential solutions to the optimization problem.
@@
Graph representation of the STSP network. The graph consists of yellow, blue, and green circles connected by black and red directed edges. The yellow circles represent the $V_{R}$ nodes, the green circle is the depot, and the blue circles are the Steiner nodes. The black edges represent the arcs of the graph, while a possible route is highlighted in red, illustrating one potential solution to the optimization problem.
@@
Table 1. Summary of the notations used in the model formulation.
Notation Description
$V$ Set of all vertices in the graph
$A$ Set of directed arcs in the graph
$V_{R}$ Set of the required nodes to visit
$\delta^{+}(i)$ Set of arcs outgoing from node $i$
$\delta^{-}(i)$ Set of arcs incoming to node $i$
$c_{k}$ Cost associated with traversing arc $k$
@@ 2.2. Mathematical Formulation
(7) $y_{k}^{t}\in\{0,1\},\quad\forall k\in A,1\leq t\leq|A|$
The objective function (1) minimizes the total travel cost across all time periods. It is expressed as the sum of the costs associated with each arc traversed, weighted by the respective cost. Constraint (2) ensures that the route starts at the depot. Constraints (3) ensure that in the first period, there are no other arcs traversed except for the one originating from the depot. Constraint (4) ensures that the sum of incoming arcs at the depot node equals the sum of outgoing arcs. Constraints (5) enforce the requirement that all required vertices must be visited at least once. Constraints (6) ensure flow conservation between consecutive time periods for each node (except the depot). They state that the number of incoming arcs to a vertex in any given period must equal the number of outgoing arcs from that node in the next period, thereby maintaining a continuous flow through the network. Constraints (7) define the domain of the decision variable. It ensures that the index $k$ belongs to the set $A$ of arcs and that the time index $t$ is defined in the interval $1\leq t\leq|A|$. The upper bound $|A|$ guarantees that the model can represent even the longest path, namely the one that traverses all arcs in the graph.
@@ 3. Solution approach
    QUBO formulation: we convert both the preprocessed and unpreprocessed versions of the STSP into QUBO formulations. The preprocessed version will be referred to as Reduced QUBO (RQUBO), while the unprocessed version will be referred to as Standard QUBO (SQUBO). The QUBO formulation is solved using two different approaches:
@@ 4. Computational study
For the instances solved using Gurobi, we have used the Gurobi ver. 11 (Gurobi Optimization, LLC, 2024) and the implementation is done entirely in Python. Also, we impose a time_limit of 10 seconds. For SILP and RILP solved with Gurobi, we consider 16 instances with increasing sizes, ranging from 4 to 19 nodes. The analysis stops beyond this threshold, as no solution is obtained within the time limit.
@@
Experimental tests are conducted within a 100x100 unit square region. The set of nodes $V$ is randomly generated by assigning each node random Cartesian coordinates within this area. An additional node, denoted as the depot node 00, is generated with random coordinates within the same region. The subset of required nodes $V_{R}\subseteq V$ is selected comprising the first $70\%$ of the total vertices in $V$, rounded down to the nearest integer. Furthermore, each arc cost $c_{k}$ is assigned a random value between 20 and 50. To facilitate the replication of this study, all instances used are publicly accessible in (Ciacco et al., 2025b).
@@
  • The third section explores the performance of QA by employing the D-Wave solver.
@@ 4.1. Scalability of PMRA using Gurobi solver
We then transform both our SILP and RILP formulations into QUBO and evaluate the impact of the PMRA in reducing the problem size. Table 3 presents a comparison of the number of variables in the SQUBO and RQUBO formulations, highlighting the effectiveness of the PMRA in reducing the number of variables across different problem sizes. It consists of the column $|V|$, representing the cardinality of the set $|V|$, followed by n.var SQUBO and n.var RQUBO, which indicate the number of variables in the SQUBO and RQUBO formulations, respectively. Finally, the GAP column shows the percentage reduction achieved. Analyzing the results, we observe that the number of variables in the SQUBO formulation grows rapidly as $|V|$ increases. When the PMRA is used, the increase in the number of variables as the size of the set V grows is more contained. In fact, a maximum reduction of 57% and an average reduction of 48% are achieved.
@@ 5. Conclusions
We utilized both the D-Wave QPU and the D-Wave LeapBQM solver in our experiments. The LeapBQM consistently outperformed the QPU-based approach, both in terms of execution time and solution quality. While the QPU demonstrated effectiveness for very small problem instances, it was generally less efficient for larger problems. The LeapBQM provided high-quality solutions and showed better scalability, even for larger problem sizes. This makes the LeapBQM a more promising and practical tool for addressing real-world STSP challenges, offering superior performance across a range of problem sizes. Furthermore, in all the tests conducted across all methods, ILP solved with Gurobi, SA and both quantum approaches (QPU and LeapBQM), we observed that the PMRA consistently enhanced the solving performance. This method proved to be effective in improving the computational efficiency and the quality of the solutions, underlining its importance as a significant contribution to solving the STSP.
For future work, a deeper exploration into why QA still provides inferior results compared to the ILP version solved by Gurobi would be highly valuable. Understanding the limitations of quantum approaches in this context could offer important insights and help identify ways to improve their performance. Additionally, further investigation into the potential of QA for solving larger and more complex instances would contribute to the development of more scalable quantum optimization techniques, making them more applicable to real-world problems. Another promising direction for future research involves exploring quadratic formulations. This approach would allow for greater flexibility in modeling and may lead to improved results, especially when handling time-dependent constraints or more intricate network structures. Finally, as the problem scope expands to larger logistics networks and more complex instances, there is potential for further refinements to the PMRA method. This would enable better handling of increasingly complicated scenarios, providing a deeper analysis of the trade-offs between classical and quantum methods.
@@ Acknowledgments
Eneko Osaba acknowledges support from the Basque Government through (EXP.2022/01341) (A/20220551) (Plan complementario de comunicación cúantica). During the preparation of this work, the authors used Microsoft Copilot to improve the language and readability of the manuscript. After using this tool/service, the authors have reviewed and edited the content as needed and take full responsibility for the content of the publication.